OpenStudy (anonymous):

Which ratio forms a proportion with 18/32?

OpenStudy (mathstudent55):

Any fraction that is equivalent to 18/32.

OpenStudy (mertsj):

\[\frac{18}{32}=\frac{36}{64}=\frac{9}{16}=\frac{54}{96}=\frac{-72}{-128}\]

OpenStudy (mertsj):

Take your pick.

OpenStudy (anonymous):

thank you so much :)

OpenStudy (mertsj):

yw